OpenStudy (anonymous):

Solve for (x+5)^2=12

OpenStudy (anonymous):

x=-5+or- 2 radical 3

OpenStudy (jamesj):

Suppose \[y^2 = 12\] What would say is the solution or are the solutions for y?

OpenStudy (amistre64):

if only we had a sqrt(12) in there but that +5 is getting in the way .... i say we put a hit out on it :) -5 +- sqrt(12)

OpenStudy (anonymous):

could it be -5-2√6,-5+2√6

OpenStudy (amistre64):

almost; but no

OpenStudy (amistre64):

2sqrt(6) = sqrt(24)

OpenStudy (anonymous):

-5-2√3,-5+2√3

OpenStudy (jamesj):

\[\sqrt{12} = \sqrt{4 \times 3} = \sqrt{2^2 \times 3} = \sqrt{2^2} \times \sqrt{3} = 2 \sqrt{3}\]

OpenStudy (anonymous):

-2√3 2√3 has to be close

OpenStudy (jamesj):

No, you were right above\[x = -5 \pm 2 \sqrt{3}\]